I would like to know the steps to factoring this problem; it is a fraction.
4x-12 over 4x

Answer by Alan3354(69443)

(4x-12)/4x
-----------
Look for even numbers first.
(4x-12)/4x = 4(x-3)/4x
= (x-3)/x
No further reduction is possible.
---------------
It can be written as 1 - (3/x), but there's no advantage to that.